OpenStudy (anonymous):

(1/r) + (r/(2-r))=-7/4 How do you solve this??

OpenStudy (anonymous):

to add fractions, their denominators must be the same. to do so we will force it by multiplying by the missing term(s). notice that were really just multiplying by one. \[\frac{ 1 }{ r } + \frac{ r }{ 2-r } = \frac{ 7 }{ 4 }\]\[\frac{ 1 }{ r }\times \frac{ 2-r }{ 2-r } + \frac{ r }{ 2-r } \times \frac{ r }{ r } = \frac{ 7 }{ 4 }\]\[\frac{ 2-r }{ r(2-r) } + \frac{ r^2 }{ r(2-r) } = \frac{ 7 }{ 4 }\]\[\frac{ 2-r + r^2 }{ r(2-r) } = \frac{ 7 }{ 4 }\]\[2-r+r^2 = \frac{ 7 }{ 4}\left( r(2-r) \right)\] and solve from here. probably best using the quadratic formula. let me know if you have follow-up questions ^_^

OpenStudy (anonymous):

Wait May you type it using the equation generator I don't really understand what youutypes D;

OpenStudy (anonymous):

i did use it. refreshing page might fix it

OpenStudy (anonymous):

mhm yea seems to not show up :\

OpenStudy (anonymous):

\[x^2 = \pi\] do you see this as proper latex? are you blocking javascript with your browser?

OpenStudy (anonymous):

hmhm yea that's what it seems

OpenStudy (anonymous):

Would writing it out be better?